OpenStudy (anonymous):

my issues are in math, I tried to do these problems on my own but I don't feel I did them right: 9z-3/8=9/16 need to solve equation

OpenStudy (anonymous):

\[\frac{9z-3}{8}=\frac{9}{16} \] \[16(9z-3) = 9 * 8\] \[144z - 48 = 72\] \[144z = 120\] so z = 120/144 = 5/6

OpenStudy (anonymous):

thank you, I had nothing like this

OpenStudy (anonymous):

yw